A toy rocket rises vertically in such a way that t seconds after its liftoff, it is
s(t) 16t 200t
2
= − +
feet above the ground.
(a) How high is the rocket after 6 seconds?
(b) What is the average velocity of the rocket over the first 6 seconds of flight (between t=0 and t=6)?
(c) What is the instantaneous velocity of the rocket at t=2 sec?
1
Expert's answer
2020-12-01T15:03:46-0500

"\displaystyle(a).\\

s(t) = -16t^2 + 200t\\

\begin{aligned}
s(6) &= -16(6)^2 + 200(6) 
\\&= 1200 - 576 = 624\,\textsf{fts}
\end{aligned}\\


(b)\\
\textsf{To determine the average velocity}\\
\textsf{of the particle between}\, 1.0 \, s\, \, \textsf{and}\,6.0 \,s,\\
\textsf{we calculate the values of}\\
s(1.0\,s) \,\, \textsf{and}\,\, s(6.0\,s):\\

\begin{aligned}
s(6) &= -16t^2 + 200t = -16(6)^2 + 200(6) \\&= 1200 - 576 = 624
\end{aligned}\\


\begin{aligned}
s(0) &= -16(0)^2 + 200(0)\\
&= 0
\end{aligned}\\

\begin{aligned}
v_{average} &= \frac{s(6) - s(0)}{6 - 0} 
\\&= \frac{624 - 0}{6 - 0} = \frac{624}{6}
\\&= 104 \,\textsf{fts/s}
\end{aligned}\\

\textbf{\textsf{OR}}\\
\textsf{Since, velocity is the rate}\\
\textsf{of change of distance with time}\\
v = \frac{\mathrm{d}(s(t))}{\mathrm{d}t} = -32t + 200\\

\begin{aligned}
v(6) &= -32(6) + 200\\
&= -192 + 200 = 8
\end{aligned}\\


\begin{aligned}
v(0) &= -32(0) + 200\\
&= 200
\end{aligned}\\

\textsf{The sum of the initial and final}\\
\textsf{velocity is divided by}\,2 \, \textsf{to find}\\
\textsf{the average.}\\
\textsf{The average velocity calculator}\\
\textsf{uses the formula that shows the}\\
\textsf{average velocity}\, (v) \, \textsf{equals}\\
\textsf{the sum of the final velocity}\, (v) \, \\
\textsf{and the initial velocity}\, (u), \, \textsf{divided by}\, 2.\\

\begin{aligned}
v_{average} &= \frac{v_6 + v_0}{2} 
\\&= \frac{200 + 8}{2} = \frac{208}{2}
\\&= 104 \,\textsf{fts/s}
\end{aligned}\\


(c)\\
\begin{aligned}
v(2) &= -32(2) + 200\\
&= -64 + 200 = 136\,\textsf{fts/s}
\end{aligned}"